Quotations from M.C. Escher’s Writings – Class on Wednesday, 3/16

With reference to his regular division of the plane drawings:

“I cannot imagine what my life would be like if this problem had never occurred to me. One might say that I am head over heels in love with it, and I still don’t know why.”

With reference to the relation of his work to the classification of the 17 types of “wallpaper symmetry groups:”

“Crystallographers have put forward a definition of the idea; they have ascertained ... how many systems or ways there are of dividing a plane in a regular manner. In so doing, they have opened the gate to an extensive domain, but they have not entered the domain themselves. By their very nature, they are more interested in the way in which the gate is opened than in the garden lying behind it. ... A long time ago I chanced upon this domain in one of my wanderings; I saw a high wall and as I had a premonition of an enigma, something that might be hidden behind the wall, I climbed over it with some difficulty. However on the other side I landed in a wilderness and had to cut my way through with a great effort until–by a circuitous route–I came to the open gate, the open gate of mathematics. From there, well-trodden paths lead in every direction and then I suddenly discover a new path and experience fresh delights ... I feel a revitalizing yet oppressive sense of loneliness. That is why I have been extolling the existence of this paradise for many years ... even though I do not expect many people to wander through. For what fascinates me ... often seems to be considered dry and tedious by others.”

With reference to one of his contacts with a mathematician (H.S.M. Coxeter – one of the great geometers of the 20th century). Note: This quotation refers to part of the genesis of the Circle Limit prints we saw last semester in connection with non-Euclidean geometries.

“His hocus-pocus text is of no use to me at all, but the picture can probably help me produce a division of the plane which promises to become an entirely new variation of my series of divisions of the plane.”

With reference to his affinity with scientists (and mathematicians):

“I have often felt closer to people who work scientifically (though I certainly do not do so myself) than to my fellow artists.”

As we said in class, his systematic explorations of colored tilings were closer to mathematical research than he might have realized(!)